# “Whose Side Are You On?” Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection

### v1

@@ 5 Conclusion
This study explored the efficacy of few-shot learning using Large Language Models for the task of ideology estimation across diverse textual datasets, including news articles and YouTube videos. Our findings demonstrate that LLMs, particularly GPT-4o, achieve state-of-the-art performance, surpassing traditional supervised baselines. By leveraging advanced few-shot in-context learning methods, we addressed challenges of scalability and robustness, showcasing the ability of LLMs to adapt to evolving contexts with minimal supervision.
@@ 7 Ethics Statement
Appendix A Prompts used for the LLM
@@ Appendix E Sources can be misleading
Our prior results showed that providing the source information significantly increases the accuracy. However, not all content from a given source may share the ideology of the source itself Haroon et al. (2023), as some liberal articles can come from conservative outlets. To highlight this particular scenario, we identify cases in our test set where the ideology of the article or video is different from the ideology of its source for all three datasets (e.g., a liberal article from Fox News or a conservative article from MSNBC). We rely on AllSides’ media bias chart555https://www.allsides.com/media-bias/media-bias-chart for source ideology and map these sources to their corresponding news articles and YouTube videos. We then compare the accuracy of GPT-4o when 1) only title is provided, 2) only source is provided, and 3) both title and source are provided.
@@
Appendix F Few-shot results for additional datasets
@@
Figures 4, 5, and 6 show the increases in accuracy across all datasets by increasing the number of shots.
Appendix G Placement of popular news channels based on the slant cut-offs
Figure 7 shows the placement of popular news channels based on the YouTube slant dataset. Corresponding cut-offs at -0.33 and +0.33 are also shown for the liberal and conservative respectively.
Appendix H Qualitative results for the political news analysis for YouTube videos
Qualitative results for political news videos, non-political news videos, and political non-news videos that changed predictions as a result of adding information about the source are provided in Tables 13, 14, and 15 respectively.
| Title | Channel |
| Liberal to Neutral | |
| How welcome are Russian deserters in Georgia? | Focus on Europe | DW News |
| Ukrainians ditch Russian traditions and celebrate Christmas with the West | Sky News |
| Asian Americans Debate Model Minority & Asian Hate | VICE Debates | VICE |
| Secret Service under scrutiny ahead January 6 committee’s final report | PBS NewsHour |
| White House Correspondent Jeff Mason on his tense exchanges with Trump | Reuters |
| Liberal to Conservative | |
| Democrats introduce bill to block Trump from holding office | Newsmax |
| LIVE: Jan. 6 Committee Votes on Criminal Referrals, Final Report | NTD |
| New CRIMINAL REFERRALS Against Former President Trump Unveiled by Jan. 6th Committee | BlazeTV |
| Tonight’s Tipping Points: Hate Crimes, Planned Parenthood, & Trans Military Members! | One America News Network |
| Trolling #AMFEST2022 Attendees as Left-Winger ’Garth Griffin’ of Edge News | Project Veritas |
| Neutral to Liberal | |
| Haberman says Trump has been diminished but isn’t dead politically | CNN |
| Kevin McCarthy’s Path To Speaker Is In Peril | MSNBC |
| Anger grows in Russia over Ukraine’s New Year’s Day strike on Russian troops | CNN |
| Rep. Adam Kinzinger Delivers Final Speech As Congressman | MSNBC |
| ’Shocked’: Russia expert explains why Putin’s change of word is significant | CNN |
| Neutral to Conservative | |
| UK Weather: What’s behind Britain’s big freeze? | UK snow chaos explained | Daily Mail |
| Transgender Clinical Psychologist on people de-transitioning | Dr. Erica Anderson | Newsmax |
| Actor slams crypto as ’biggest Ponzi scheme ever’ during Senate hearing | Fox News |
| Rep. Spartz explains her reasoning for voting ’present’ for Speaker | Fox News |
| Internet users allege Bryan Kohberger attended Idaho victims’ vigil | Fox News |
| Conservative to Liberal | |
| Trump Supporter: Too Much Diversity Will Kill America | The Young Turks |
| Russian foreign minister: Ukraine must give up occupied territories | CNN |
| These Trump Supporters Are CONVINCED "Sleepy Joe Poopy Pants Isn’t The Acting President" | The Young Turks |
| Steve Bannon reacts to Trump’s ’major announcement’ | CNN |
| Trump adviser tells reporter ’the magic is gone’ in new article about re-election bid | CNN |
| Conservative to Neutral | |
| Justin Amash APPLAUDS Revolt Against McCarthy, Says Ruling Oligarchy CRUSHES Dissent | The Hill |
| State Of Emergency Declared At Texas Border | NBC News |
| Watch Trump claim the army retook airports during the Revolutionary War in July Fourth speech | The Globe and Mail |
| Trump Calls Jan. 6 Committee A ’Kangaroo Court’ After Criminal Referrals | NBC News |
| ’They shot up the WRONG house!’ Sheriff Grady Judd on young kids shot at while in bed in Polk County | WFLA News Channel 8 |
| Title | Channel |
| Liberal to Neutral | |
| TRIPLEDEMIC Or Free Market Failure To Blame For Medicine SHORTAGE? STRUGGLE To Find Children’s Meds | The Hill |
| Watch: Jan. 6 Committee Presents Key Video Evidence Compilation | NBC News |
| Neutral to Liberal | |
| Top U.S. & World Headlines — December 27, 2022 | Democracy Now! |
| Top U.S. & World Headlines — January 3, 2023 | Democracy Now! |
| Raskin On Cancer Diagnosis: ’Totally My Plan To Make It Through This Thing’ | MSNBC |
| Watch The 11th Hour Highlights: Dec. 20 | MSNBC |
| Plans For Zelenskyy D.C. Trip Came Together In A ’Matter Of Days’ | MSNBC |
| Neutral to Conservative | |
| McCarthy speakership battle: What happens if he concedes? | Fox Business |
| The yule log Christmas tradition | Fox News |
| “Woman” Is Named Word of the Year | DailyWire+ |
| What is the controversy surrounding Meghan Markle and her sister? | Fox News |
| ’DOUBLE WHAMMY’: Real estate expert issues pricing warning | Fox Business |
| Conservative to Liberal | |
| All Gestures, All Shock Opera: DeSantis Launches His Newest Battle | MSNBC |
| Christian Influencer Makes Surprising Confession About Trans Porn | The Young Turks |
| The War on Christmas RAGES ON! | The Young Turks |
| Alex Jones Says He’s Paying His Personal Trainer $100,000 A Week | The Young Turks |
| ’Look at where he is today’: Bash on Kinzinger’s farewell speech | CNN |
| Conservative to Neutral | |
| Cancelling the F-22 Fighter is Worse than You Think | Task & Purpose |
| Title | Channel |
| Liberal to Neutral | |
| Is capitalism actually broken? | TED-Ed |
| Update from Ukraine | Ukraine is getting ready for the winter Attack | Ruzzia will lose it | Denys Davydov |
| Van Lathan: A Jewish Producer Confronted Kanye at TMZ when Ye Said He Loves Hitler (Part 2) | djvlad |
| The Trouble with the Electoral College | CGP Grey |
| Those Times Sam Harris Made Us Slap Our Knees | Enlightainment |
| Liberal to Conservative | |
| The New Beginning of Same-Sex Marriage | Ben Shapiro |
| Meet Kamala Harris’s 6-Year-Old Speechwriter | The Babylon Bee |
| Pro-Choice Bosses - A Generous Offer To Our Pregnant Employees | Live Action |
| Confronting Charlie Kirk and Tim Pool at Turning Point USA America Fest | Alex Stein |
| I Showed Up To An Anti-Gun Protest In Los Angeles | Let’s Talk | James Klug |
| Neutral to Liberal | |
| Meanwhile… Pantone’s Color of the Year | Macron Makes Condoms Free In France | The Late Show with Stephen Colbert |
| Grandma Arrested For Feeding The Homeless?! Danbury Court Case Clarifcation + GIVEAWAY! | Long Island Audit |
| LIVE: EMERGENCY January 6 Committee hearing - Day 6 | Brian Tyler Cohen |
| The Funniest Donald Trump Stories On The Graham Norton Show | Part One | The Graham Norton Show |
| Prison Officer BUSTED After Massive Nazi Gang Attack Caught | The Damage Report |
| Neutral to Conservative | |
| Winning With The Hand You’re Dealt with Rick Warren | Saddleback Church |
| Fighting the Darkest Criminal Enterprise on the Planet | The 700 Club |
| STORE WARS: Target, Walmart & Home Depot go to war against Organized Retail theft | News For Reasonable People |
| Warner MAY BE SUED over Amber Heard and Aquaman 2?! NEW Hollywood Lawsuit! | ThatUmbrellaGuy |
| Historical alteration on the Ukraine Map: Russian 1st Corps withdrawing en masse from 4 locations! | Divine Justice |
| Conservative to Liberal | |
| Elon Musk Finally CRACKS In Ultimate Outburst Over Free Speech | The Damage Report |
| Chris Wallace: This Would Never Have Happened Under Nancy Pelosi | The Late Show with Stephen Colbert |
| Mike Lindell Snaps After Hannity Admits There Was No Fraud | Tommy Campbell |
| Trump ATTACKS the FBI and DOJ and Loses his Mind in NEW POSTS | MeidasTouch |
| BREAKING: Kari Lake slammed with brutal legal news | Brian Tyler Cohen |
| Conservative to Neutral | |
| Why US Let Russia Become the Biggest Nuclear Power | The Infographics Show |
| The CIA’s Biggest Scam Exposed | KiraTV |
| Cop Pulls Over "Legal Expert" And SCHOOLS HIM On The Law | Audit the Audit |
| Why US Enemies are Scared of Ghosts (Elite Special Forces) | The Infographics Show |
| "INSPIRATIONAL HOLIDAY VIDEO" — A Bad Lip Reading of Joe Biden | Bad Lip Reading |
Appendix I Code and Reproducibility

### v2

@@ 6 Conclusion
We explored the efficacy of few-shot learning using LLMs for the task of ideology estimation across diverse textual datasets, including news articles and YouTube videos. Our findings demonstrate that LLMs, particularly GPT-4o, achieve state-of-the-art performance in the few-shot setting, surpassing traditional supervised baselines. By leveraging set-cover based ICL, we addressed challenges of scalability and robustness, showcasing the ability of LLMs to adapt to evolving contexts with minimal supervision.
@@ 8 Ethics Statement
Appendix A Prompts used for LLMs
@@ Appendix E Sources can be misleading
Our prior results showed that providing the source information significantly increases the accuracy. However, not all content from a given source may share the ideology of the source itself Haroon et al. (2023), as some liberal articles can come from conservative outlets. To highlight this particular scenario, we identify cases in our test set where the ideology of the article or video is different from the ideology of its source for all three datasets (e.g., a liberal article from Fox News or a conservative article from MSNBC). We rely on AllSides’ media bias chart222https://www.allsides.com/media-bias/media-bias-chart for source ideology and map these sources to their corresponding news articles and YouTube videos. We then compare the accuracy of GPT-4o when 1) only title is provided, 2) only source is provided, and 3) both title and source are provided.
@@
Appendix F Statistical significances
We use the McNemar test to determine the statistical differences between the accuracies of two specific classifiers. Tables 10, 11, and 12 show the significance values for the models trained on the YouTube Slant, Ad Fontes news, and AllSides news datasets respectively.
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 0.13 | 0.06 | 33.06** | 0.05 | 23.58** | 8.77** | 34.18** | 0.21 | 11.05** | 10.75** | 2.56 | 39.32** | 62.01** | 4.50* | 4.21* | 2.34 | 0.39 | 1.46 | 46.89** | 15.42** |
| TS | 0.13 | — | 0.52 | 27.95** | 0.52 | 30.99** | 13.77** | 35.79** | 0.69 | 12.10** | 12.30** | 4.69* | 41.32** | 69.72** | 4.72* | 2.79 | 3.31 | 0.04 | 2.77 | 53.35** | 20.66** | |
| TB | 0.06 | 0.52 | — | 42.29** | 0.00 | 28.96** | 10.27** | 30.30** | 0.06 | 10.25** | 10.16** | 2.13 | 37.80** | 65.91** | 2.96 | 5.33* | 1.70 | 0.88 | 1.07 | 54.30** | 18.06** | |
| TSB | 33.06** | 27.95** | 42.29** | — | 25.95** | 89.19** | 67.10** | 89.28** | 22.45** | 59.74** | 58.72** | 33.34** | 96.75** | 132.08** | 41.44** | 7.87** | 45.75** | 22.90** | 29.36** | 119.71** | 72.90** | |
| S | 0.05 | 0.52 | 0.00 | 25.95** | — | 18.40** | 8.13** | 28.34** | 0.03 | 7.84** | 8.07** | 3.54 | 34.04** | 65.24** | 2.02 | 5.47* | 0.98 | 0.76 | 1.90 | 38.75** | 12.04** | |
| B | 23.58** | 30.99** | 28.96** | 89.19** | 18.40** | — | 4.99* | 3.06 | 13.20** | 0.96 | 0.85 | 5.33* | 5.84* | 20.22** | 5.08* | 42.11** | 11.61** | 29.94** | 6.74** | 7.38** | 0.47 | |
| SB | 8.77** | 13.77** | 10.27** | 67.10** | 8.13** | 4.99* | — | 10.69** | 4.56* | 0.25 | 0.30 | 0.73 | 15.08** | 38.23** | 0.48 | 23.46** | 2.41 | 14.00** | 1.43 | 20.55** | 1.47 | |
| Llama | T | 34.18** | 35.79** | 30.30** | 89.28** | 28.34** | 3.06 | 10.69** | — | 37.09** | 9.03** | 8.62** | 16.12** | 0.35 | 6.28* | 19.19** | 58.08** | 23.16** | 43.91** | 18.64** | 0.02 | 5.61* |
| TS | 0.21 | 0.69 | 0.06 | 22.45** | 0.03 | 13.20** | 4.56* | 37.09** | — | 7.81** | 8.57** | 1.61 | 35.74** | 59.09** | 1.75 | 5.46* | 0.56 | 1.20 | 0.64 | 32.53** | 9.71** | |
| TB | 11.05** | 12.10** | 10.25** | 59.74** | 7.84** | 0.96 | 0.25 | 9.03** | 7.81** | — | 0.00 | 1.69 | 15.84** | 36.84** | 1.71 | 26.06** | 4.14* | 16.75** | 2.70 | 10.61** | 0.16 | |
| TSB | 10.75** | 12.30** | 10.16** | 58.72** | 8.07** | 0.85 | 0.30 | 8.62** | 8.57** | 0.00 | — | 1.92 | 15.09** | 35.27** | 1.83 | 26.38** | 4.38* | 17.32** | 2.98 | 9.67** | 0.11 | |
| S | 2.56 | 4.69* | 2.13 | 33.34** | 3.54 | 5.33* | 0.73 | 16.12** | 1.61 | 1.69 | 1.92 | — | 23.06** | 45.22** | 0.00 | 14.49** | 0.18 | 5.32* | 0.20 | 17.86** | 3.01 | |
| B | 39.32** | 41.32** | 37.80** | 96.75** | 34.04** | 5.84* | 15.08** | 0.35 | 35.74** | 15.84** | 15.09** | 23.06** | — | 5.76* | 21.04** | 69.73** | 31.29** | 53.60** | 25.26** | 0.14 | 10.29** | |
| SB | 62.01** | 69.72** | 65.91** | 132.08** | 65.24** | 20.22** | 38.23** | 6.28* | 59.09** | 36.84** | 35.27** | 45.22** | 5.76* | — | 38.22** | 92.66** | 51.61** | 76.55** | 48.80** | 6.19* | 30.08** | |
| Mistral | T | 4.50* | 4.72* | 2.96 | 41.44** | 2.02 | 5.08* | 0.48 | 19.19** | 1.75 | 1.71 | 1.83 | 0.00 | 21.04** | 38.22** | — | 18.21** | 0.43 | 7.90** | 0.17 | 19.38** | 2.88 |
| TS | 4.21* | 2.79 | 5.33* | 7.87** | 5.47* | 42.11** | 23.46** | 58.08** | 5.46* | 26.06** | 26.38** | 14.49** | 69.73** | 92.66** | 18.21** | — | 13.56** | 2.79 | 11.83** | 69.98** | 32.90** | |
| TB | 2.34 | 3.31 | 1.70 | 45.75** | 0.98 | 11.61** | 2.41 | 23.16** | 0.56 | 4.14* | 4.38* | 0.18 | 31.29** | 51.61** | 0.43 | 13.56** | — | 9.75** | 0.00 | 43.68** | 8.44** | |
| TSB | 0.39 | 0.04 | 0.88 | 22.90** | 0.76 | 29.94** | 14.00** | 43.91** | 1.20 | 16.75** | 17.32** | 5.32* | 53.60** | 76.55** | 7.90** | 2.79 | 9.75** | — | 3.62 | 68.73** | 24.89** | |
| S | 1.46 | 2.77 | 1.07 | 29.36** | 1.90 | 6.74** | 1.43 | 18.64** | 0.64 | 2.70 | 2.98 | 0.20 | 25.26** | 48.80** | 0.17 | 11.83** | 0.00 | 3.62 | — | 20.30** | 4.24* | |
| B | 46.89** | 53.35** | 54.30** | 119.71** | 38.75** | 7.38** | 20.55** | 0.02 | 32.53** | 10.61** | 9.67** | 17.86** | 0.14 | 6.19* | 19.38** | 69.98** | 43.68** | 68.73** | 20.30** | — | 20.33** | |
| SB | 15.42** | 20.66** | 18.06** | 72.90** | 12.04** | 0.47 | 1.47 | 5.61* | 9.71** | 0.16 | 0.11 | 3.01 | 10.29** | 30.08** | 2.88 | 32.90** | 8.44** | 24.89** | 4.24* | 20.33** | — | |
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 34.40** | 20.11** | 449.51** | 341.98** | 188.99** | 377.91** | 96.77** | 32.68** | 38.99** | 224.78** | 66.33** | 47.50** | 172.07** | 51.50** | 2.29 | 18.13** | 288.68** | 232.80** | 135.35** | 287.96** |
| TS | 34.40** | — | 6.67** | 435.86** | 328.99** | 158.38** | 362.25** | 200.31** | 132.25** | 140.91** | 218.99** | 60.72** | 42.21** | 163.93** | 136.48** | 19.91** | 82.16** | 265.40** | 231.29** | 113.12** | 264.55** | |
| TB | 20.11** | 6.67** | — | 440.16** | 344.31** | 173.22** | 374.56** | 163.22** | 81.73** | 103.88** | 233.55** | 68.50** | 50.77** | 176.20** | 104.43** | 3.12 | 62.26** | 283.78** | 240.87** | 128.75** | 283.02** | |
| TSB | 259.15** | 373.77** | 318.02** | — | 24.89** | 175.03** | 15.45** | 68.03** | 125.03** | 124.95** | 107.43** | 249.53** | 286.51** | 146.23** | 130.00** | 286.31** | 175.39** | 83.12** | 75.74** | 204.39** | 74.29** | |
| S | 253.85** | 372.88** | 325.06** | 24.89** | — | 61.13** | 4.88* | 73.52** | 136.62** | 132.00** | 36.33** | 186.29** | 197.09** | 72.16** | 134.29** | 293.99** | 177.19** | 12.26** | 33.39** | 93.04** | 14.84** | |
| B | 199.10** | 272.44** | 239.34** | 175.03** | 61.13** | — | 91.51** | 28.50** | 61.03** | 65.13** | 7.92** | 20.40** | 40.57** | 0.19 | 75.61** | 192.91** | 120.69** | 33.56** | 13.44** | 8.85** | 33.00** | |
| SB | 245.87** | 363.62** | 310.64** | 15.45** | 4.88* | 91.51** | — | 64.08** | 121.96** | 118.26** | 51.81** | 201.48** | 222.84** | 96.82** | 127.00** | 283.93** | 168.95** | 24.57** | 44.52** | 124.66** | 31.57** | |
| Llama | T | 96.77** | 200.31** | 163.22** | 433.77** | 350.42** | 149.02** | 375.46** | — | 21.99** | 21.16** | 236.90** | 72.80** | 54.29** | 180.28** | 18.25** | 138.91** | 40.15** | 276.51** | 263.02** | 114.96** | 266.27** |
| TS | 32.68** | 132.25** | 81.73** | 475.86** | 395.04** | 164.84** | 413.03** | 21.99** | — | 0.00 | 274.00** | 103.51** | 75.98** | 215.70** | 0.00 | 85.23** | 5.80* | 299.13** | 317.34** | 131.29** | 295.40** | |
| TB | 38.99** | 140.91** | 103.88** | 468.49** | 379.28** | 163.60** | 400.96** | 21.16** | 0.00 | — | 271.24** | 99.85** | 76.18** | 220.90** | 0.00 | 81.82** | 8.23** | 299.39** | 307.37** | 133.30** | 297.66** | |
| TSB | 252.24** | 383.38** | 327.86** | 107.43** | 36.33** | 7.92** | 51.81** | 79.93** | 148.70** | 144.11** | — | 78.32** | 130.33** | 12.27** | 136.57** | 313.88** | 184.20** | 7.11** | 2.32 | 28.24** | 5.32* | |
| S | 262.93** | 399.75** | 341.51** | 249.53** | 186.29** | 20.40** | 201.48** | 96.50** | 180.50** | 174.87** | 78.32** | — | 2.49 | 46.08** | 160.35** | 341.84** | 201.12** | 102.45** | 121.28** | 7.80** | 103.54** | |
| B | 260.62** | 393.63** | 345.41** | 286.51** | 197.09** | 40.57** | 222.84** | 107.51** | 190.68** | 185.48** | 130.33** | 2.49 | — | 88.67** | 164.74** | 343.47** | 215.15** | 147.84** | 126.48** | 23.21** | 135.88** | |
| SB | 265.39** | 390.41** | 337.63** | 146.23** | 72.16** | 0.19 | 96.82** | 89.64** | 161.96** | 160.99** | 12.27** | 46.08** | 88.67** | — | 153.62** | 325.90** | 201.51** | 27.24** | 18.31** | 8.15** | 27.36** | |
| Mistral | T | 51.50** | 136.48** | 104.43** | 451.99** | 360.36** | 167.31** | 400.83** | 18.25** | 0.00 | 0.00 | 237.93** | 79.88** | 57.59** | 186.23** | — | 86.05** | 9.88** | 282.52** | 272.88** | 131.00** | 295.80** |
| TS | 2.29 | 19.91** | 3.12 | 465.55** | 366.27** | 173.09** | 401.68** | 138.91** | 85.23** | 81.82** | 249.35** | 84.57** | 60.26** | 194.64** | 86.05** | — | 33.00** | 295.92** | 276.45** | 131.27** | 291.16** | |
| TB | 18.13** | 82.16** | 62.26** | 453.61** | 360.22** | 187.09** | 393.63** | 40.15** | 5.80* | 8.23** | 240.67** | 76.52** | 59.10** | 191.91** | 9.88** | 33.00** | — | 295.82** | 255.13** | 143.73** | 299.24** | |
| TSB | 240.52** | 343.16** | 295.33** | 83.12** | 12.26** | 33.56** | 24.57** | 56.78** | 104.94** | 107.75** | 7.11** | 102.45** | 147.84** | 27.24** | 113.84** | 263.56** | 161.16** | — | 0.71 | 74.25** | 0.02 | |
| S | 270.72** | 411.28** | 350.63** | 75.74** | 33.39** | 13.44** | 44.52** | 97.52** | 180.09** | 169.62** | 2.32 | 121.28** | 126.48** | 18.31** | 163.77** | 345.38** | 205.33** | 0.71 | — | 32.63** | 0.51 | |
| B | 195.26** | 276.06** | 247.61** | 204.39** | 93.04** | 8.85** | 124.66** | 29.48** | 64.62** | 69.14** | 28.24** | 7.80** | 23.21** | 8.15** | 77.65** | 203.68** | 125.25** | 74.25** | 32.63** | — | 77.30** | |
| SB | 219.47** | 317.90** | 275.13** | 74.29** | 14.84** | 33.00** | 31.57** | 47.58** | 94.93** | 95.66** | 5.32* | 103.54** | 135.88** | 27.36** | 104.22** | 243.56** | 146.45** | 0.02 | 0.51 | 77.30** | — | |
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 277.85** | 32.27** | 255.80** | 437.95** | 16.75** | 383.38** | 23.89** | 60.47** | 6.33* | 25.05** | 280.54** | 48.25** | 141.66** | 17.31** | 186.27** | 47.23** | 33.34** | 339.27** | 106.59** | 99.26** |
| TS | 277.85** | — | 195.62** | 3.90* | 139.76** | 316.95** | 94.82** | 258.18** | 26.08** | 212.25** | 81.25** | 28.02** | 301.64** | 3.96* | 272.09** | 4.23* | 357.08** | 139.18** | 62.00** | 448.55** | 43.41** | |
| TB | 32.27** | 195.62** | — | 192.83** | 359.77** | 74.74** | 311.69** | 56.81** | 26.50** | 30.13** | 3.88* | 212.64** | 91.74** | 84.88** | 52.94** | 118.53** | 107.43** | 4.20* | 262.49** | 184.46** | 49.97** | |
| TSB | 255.80** | 3.90* | 192.83** | — | 150.75** | 318.84** | 109.80** | 240.60** | 17.80** | 200.62** | 70.24** | 38.68** | 290.01** | 0.96 | 253.02** | 0.56 | 342.79** | 124.91** | 73.86** | 437.54** | 31.74** | |
| S | 437.95** | 139.76** | 359.77** | 150.75** | — | 518.92** | 27.23** | 483.85** | 187.55** | 434.56** | 284.06** | 20.53** | 540.53** | 150.78** | 481.42** | 149.24** | 547.39** | 324.18** | 57.02** | 656.50** | 263.31** | |
| B | 16.75** | 316.95** | 74.74** | 318.84** | 518.92** | — | 480.17** | 2.13 | 97.45** | 0.63 | 57.42** | 381.03** | 15.78** | 215.47** | 0.13 | 239.73** | 6.63* | 81.30** | 425.04** | 67.27** | 202.21** | |
| SB | 383.38** | 94.82** | 311.69** | 109.80** | 27.23** | 480.17** | — | 441.35** | 150.71** | 393.49** | 243.99** | 7.00** | 508.99** | 119.71** | 437.53** | 109.73** | 505.97** | 279.39** | 21.33** | 619.11** | 229.59** | |
| Llama | T | 23.89** | 258.18** | 56.81** | 240.60** | 483.85** | 2.13 | 441.35** | — | 247.06** | 12.31** | 132.97** | 427.63** | 7.13** | 249.42** | 2.17 | 260.91** | 0.13 | 80.77** | 412.95** | 14.41** | 145.35** |
| TS | 60.47** | 26.08** | 26.50** | 17.80** | 187.55** | 97.45** | 150.71** | 247.06** | — | 164.01** | 24.63** | 106.58** | 224.01** | 12.95** | 162.56** | 17.64** | 151.72** | 16.56** | 119.78** | 203.03** | 0.07 | |
| TB | 6.33* | 212.25** | 30.13** | 200.62** | 434.56** | 0.63 | 393.49** | 12.31** | 164.01** | — | 140.34** | 356.14** | 39.17** | 223.88** | 2.27 | 204.32** | 10.70** | 50.39** | 357.48** | 41.93** | 109.47** | |
| TSB | 25.05** | 81.25** | 3.88* | 70.24** | 284.06** | 57.42** | 243.99** | 132.97** | 24.63** | 140.34** | — | 192.01** | 179.36** | 75.61** | 89.82** | 71.34** | 108.30** | 0.17 | 205.99** | 159.25** | 17.00** | |
| S | 280.54** | 28.02** | 212.64** | 38.68** | 20.53** | 381.03** | 7.00** | 427.63** | 106.58** | 356.14** | 192.01** | — | 510.90** | 72.76** | 365.40** | 51.85** | 441.71** | 207.56** | 1.19 | 541.63** | 131.80** | |
| B | 48.25** | 301.64** | 91.74** | 290.01** | 540.53** | 15.78** | 508.99** | 7.13** | 224.01** | 39.17** | 179.36** | 510.90** | — | 368.67** | 14.31** | 285.29** | 4.71* | 123.77** | 474.60** | 3.24 | 211.43** | |
| SB | 141.66** | 3.96* | 84.88** | 0.96 | 150.78** | 215.47** | 119.71** | 249.42** | 12.95** | 223.88** | 75.61** | 72.76** | 368.67** | — | 214.00** | 0.12 | 281.15** | 69.76** | 93.24** | 372.05** | 15.48** | |
| Mistral | T | 17.31** | 272.09** | 52.94** | 253.02** | 481.42** | 0.13 | 437.53** | 2.17 | 162.56** | 2.27 | 89.82** | 365.40** | 14.31** | 214.00** | — | 292.91** | 4.31* | 87.31** | 396.10** | 30.63** | 140.76** |
| TS | 186.27** | 4.23* | 118.53** | 0.56 | 149.24** | 239.73** | 109.73** | 260.91** | 17.64** | 204.32** | 71.34** | 51.85** | 285.29** | 0.12 | 292.91** | — | 339.67** | 138.56** | 77.25** | 391.42** | 20.98** | |
| TB | 47.23** | 357.08** | 107.43** | 342.79** | 547.39** | 6.63* | 505.97** | 0.13 | 151.72** | 10.70** | 108.30** | 441.71** | 4.71* | 281.15** | 4.31* | 339.67** | — | 175.10** | 462.37** | 26.37** | 227.88** | |
| TSB | 33.34** | 139.18** | 4.20* | 124.91** | 324.18** | 81.30** | 279.39** | 80.77** | 16.56** | 50.39** | 0.17 | 207.56** | 123.77** | 69.76** | 87.31** | 138.56** | 175.10** | — | 232.36** | 226.26** | 35.97** | |
| S | 339.27** | 62.00** | 262.49** | 73.86** | 57.02** | 425.04** | 21.33** | 412.95** | 119.78** | 357.48** | 205.99** | 1.19 | 474.60** | 93.24** | 396.10** | 77.25** | 462.37** | 232.36** | — | 569.29** | 174.80** | |
| B | 106.59** | 448.55** | 184.46** | 437.54** | 656.50** | 67.27** | 619.11** | 14.41** | 203.03** | 41.93** | 159.25** | 541.63** | 3.24 | 372.05** | 30.63** | 391.42** | 26.37** | 226.26** | 569.29** | — | 365.85** | |
| SB | 99.26** | 43.41** | 49.97** | 31.74** | 263.31** | 202.21** | 229.59** | 145.35** | 0.07 | 109.47** | 17.00** | 131.80** | 211.43** | 15.48** | 140.76** | 20.98** | 227.88** | 35.97** | 174.80** | 365.85** | — | |
Appendix G Few-shot results for additional datasets
@@
Figures 3, 4, and 5 show the increases in accuracy across all datasets by increasing the number of shots.
Appendix H Code and Reproducibility